OpenStudy (anonymous):

Convert r=12 cos(theta) to rectangular form

satellite73 (satellite73):

\[x=r\cos(\theta)\\ \frac{x}{r}=\cos(\theta)\]

satellite73 (satellite73):

substitute get \[\frac{r}{12}=\frac{x}{r}\]

satellite73 (satellite73):

making \[r^2=12x\] or \[x^2+y^2=12x\]

OpenStudy (anonymous):

I got to x^2+y^2=12x, but wasn't sure if that is rectangular form. Can you confirm that it is?

satellite73 (satellite73):

you could put \[x^2+y^2-12x=0\] or complete the square to make it a circle in standard form

OpenStudy (anonymous):

Thanks for the help!